# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

### v2

@@ Abstract.
The release of ChatGPT in 2022 triggered a rapid surge in generative artificial intelligence mobile apps (i.e., Gen-AI apps). Despite widespread adoption, little is known about how end users perceive and evaluate these Gen-AI functionalities in practice. In this work, we conduct a user-centered analysis of 676,066 reviews from 173 Gen-AI apps on the Google Play Store. We introduce a four-phase methodology, SARA (Selection, Acquisition, Refinement, and Analysis), that enables the systematic extraction of user insights using prompt-based LLM techniques. First, we demonstrate the reliability of LLMs in topic extraction, achieving 91% accuracy through five-shot prompting and non-informative review filtering. Then, we apply this method to the informative reviews, identify the top 10 user-discussed topics (e.g., AI Performance, Content Quality, and Content Policy & Censorship) and analyze the key challenges and emerging opportunities. Finally, we examine how these topics evolve over time, offering insight into shifting user expectations and engagement patterns with Gen-AI apps. Based on our findings and observations, we present actionable implications for developers and researchers.
@@
ccs: Software and its engineering Software maintenance tools
@@ 1. Introduction
Mobile applications (apps) are a crucial aspect of people’s lives, and their global data usage continues to rise each year (Rathod and Agal, 2023). A major shift in the mobile app landscape occurred with the release of ChatGPT [^1] https://play.google.com/store/apps/details?id=com.openai.chatgpt&hl=en, which achieved 7.4 million downloads within its first ten days. This rapid adoption signalled strong consumer interest in generative artificial intelligence (Gen-AI) and sparked a surge of similar artificial intelligence (AI) powered apps across app marketplaces. In 2024, this trend contributed to a 112% year-over-year increase in downloads of AI chatbot apps (Ceci, 2025), such as Microsoft Copilot [^2] https://play.google.com/store/apps/details?id=com.microsoft.copilot and DeepSeek [^3] https://play.google.com/store/apps/details?id=com.deepseek.chat. Gen-AI refers to a class of AI that generates new and meaningful content, including text, images, audio and video (Feuerriegel et al., 2024; Zhang et al., 2024). Unlike traditional AI models, which rely on historical data to perform predictions and analytical tasks, Gen-AI models synthesize information from diverse sources to create original outputs (Gozalo-Brizuela and Garrido-Merchán, 2023).
@@
Among the most prominent forms of Gen-AI are large language models (LLMs), i.e., deep learning models trained on massive text corpora to understand and generate human-like language (Hadi et al., 2023). LLMs power a wide range of natural language processing (NLP) tasks, including text generation, summarization, translation, and dialogue. Developers are increasingly integrating Gen-AI models into mobile apps to enable novel capabilities, such as conversational interfaces, voice interaction, and content generation (Hau et al., 2025). When embedded into apps, these models can deliver personalized and adaptive experiences by providing real-time, context-aware feedback that enhances user engagement and motivation (Yuen and Schlote, 2024). A key enabler of this integration is the growing availability of Gen-AI-powered application programming interfaces (APIs). APIs serve as software intermediaries that enable mobile apps to access language understanding and generation functionalities without having to host or fine-tune models locally (Chen et al., 2025b).
@@
While the integration of Gen-AI into mobile apps offers clear benefits in functionality and user experience, it also introduces a range of technical and practical challenges. Implementing Gen-AI features requires consideration of API integration, prompt engineering, token limitation, inference cost, and the unpredictable model outputs (Chen et al., 2025a). These issues can degrade the quality of Gen-AI-powered apps, leading to functional failures (e.g., crashes, incorrect behaviour, or suboptimal user interfaces), reduced efficiency (e.g., latency or excessive token consumption), and potential security vulnerabilities (Shao et al., 2025). Additionally, beyond implementation, quality assurance for Gen-AI-powered apps remains a concern. Unlike traditional software systems, Gen-AI components often lack formal specifications, making it challenging to evaluate correctness or performance in a deterministic way (Nahar et al., 2024). As a result, ensuring the reliability and robustness of Gen-AI-integrated mobile apps continues to be a significant challenge for developers.
@@
However, despite growing interest in integrating LLMs into mobile apps, little is known about the specific concerns and experiences of end-users. Existing research primarily focused on the challenges faced by developers. For example, Nahar et al. (Nahar et al., 2024) employed interviews and surveys to investigate how developers address Gen-AI integration issues, including latency, energy consumption, cost, and fairness. While informative, their work is limited to developer perspectives and does not consider how these technical challenges impact end-user experience. Chen et al. (Chen et al., 2025a) analyzed posts from an OpenAI developer forum to uncover challenges in prompting, API use, and plugin development. This study offers insight into implementation bottlenecks but also neglects user-facing concerns. Similarly, Shao et al. (Shao et al., 2025) examined open-source mobile apps with Gen-AI integration and found that 77% experienced integration challenges, which impacted functionality, performance, or security issues. However, their analysis is limited to code-level integration problems and does not investigate whether users are aware of or affected by these issues.
@@
In contrast, our work examines the user-perceived impact of Gen-AI-enabled apps, using large-scale review data to understand how users perceive Gen-AI features in practice. Existing studies on user perspectives of Gen-AI apps tend to focus on specific domains, such as education (Golding et al., 2024; Kim et al., 2025b; Lee et al., 2024a; Shata and Hartley, 2025; Kim et al., 2025a; Yuen and Schlote, 2024), creative image generation (Tang et al., 2024; Haase et al., 2023; Oppenlaender et al., 2023), or mental health (Jin et al., 2025; Heinz et al., 2025). Our work takes a systematic view by analyzing user reviews across a wide range of Gen-AI tools and application areas. To our knowledge, this is the first study to systematically examine the full landscape of topics users discuss across diverse Gen-AI domains. We identify the technical pain points surfaced by users, in addition to emerging behavioural patterns, expectations, and sources of satisfaction or frustration. This user-centered perspective complements developer-focused studies, providing critical insights for improving the usability and effectiveness of Gen-AI mobile apps.
@@
To gain these insights, we analyze user reviews that offer valuable perspectives on the challenges, expectations, and satisfaction associated with Gen-AI apps. For example, Figure 1 illustrates representative reviews from three Gen-AI apps, highlighting both praise and frustration. Building on this foundation, we introduce a four-phase methodology, Selection, Acquisition, Refinement, and Analysis (SARA). Our proposed methodology enables the large-scale collection and processing of user reviews from Gen-AI apps, supporting effective analysis using LLMs. First, we identify relevant Gen-AI apps (i.e., Selection) and collect user reviews from the Google Play Store (i.e., Acquisition). We then apply a data cleaning pipeline to enhance the quality of input data for LLM-based analysis (i.e., Refinement). Finally, we employ prompt-based LLM techniques to analyze the reviews (i.e., Analysis). Using this methodology, we conduct an empirical study on 676,066 reviews from 173 Gen-AI apps to address the following research questions (RQs):
@@
{adjustwidth}
0.8cm In this RQ, we investigate how prompt configuration (e.g., using 0-shot, 3-shot, and 5-shot) and input quality (e.g., the proportion of informative reviews) affect the performance of LLMs in extracting topics from user reviews. Our results show that increasing the number of few-shot examples and filtering non-informative reviews significantly improve accuracy. The best-performing setup achieves an overall accuracy of 91%, confirming the reliability of our LLM pipeline for downstream analysis.
RQ2: What are the most prominent topics discussed in user reviews of Gen-AI apps?
{adjustwidth}
0.8cm To answer this RQ, we apply the LLM-based topic labelling pipeline from RQ1 to informative reviews across 173 Gen-AI apps. We analyze the top 10 topics in Gen-AI app reviews to understand user behaviour, key challenges and opportunities, and derive actionable insights for developers and researchers. Our findings reveal that users tend to rate Gen-AI topics positively and expect high performance, often providing constructive feedback from a collaborative mindset. Notably, some users form emotional bonds with conversational agents and value personalized interactions. These insights have practical implications for developers aiming to improve the usability, engagement, and trustworthiness of Gen-AI features.
RQ3: What temporal trends emerge in user feedback on Gen-AI topics? {adjustwidth}0.8cm In this RQ, we explore how user feedback on Gen-AI topics evolves as the technology matures. We cluster topics to identify evolutionary trends and support the analysis by qualitative review interpretation. Our analysis reveals that while the average rating for Gen-AI topics remains high, topic trends diverge: Content Quality ratings decline due to rising user expectations, AI Performance remains stable due to balancing effects, and Content Policy & Censorship ratings improve with better filtering mechanisms with time. These shifts reflect an increasingly critical and experienced user base that raises expectations for Gen-AI apps over time.
@@
  • We propose a four-phase methodology, SARA, for analyzing user reviews of Gen-AI apps. Applying this methodology, we conduct a large-scale empirical study of 676,066 reviews from 173 Gen-AI apps on the Google Play Store to 1) identify the most frequently discussed Gen-AI-related topics, 2) uncover user expectations and pain points, and 3) analyze temporal trends to reveal how user perceptions evolve.
  • We evaluate the effectiveness of LLMs for topic extraction and assignment of user reviews, demonstrating that increasing the number of labelled examples and filtering out non-informative reviews significantly improves accuracy.
@@
  • We provide a replication package that includes the dataset of Gen-AI apps, three manually labelled samples for evaluating the filtering prompt, and the complete set of code we used to run all prompt-based experiments presented in this paper.
We organize the remainder of this paper as follows. Section 2 illustrates existing work on analyzing user reviews. Section 3 presents our methodology. Sections 46 describe the approach and the results obtained in our RQs. Section 7 synthesizes our key findings and implications of our work. Section 8 discusses potential threats to the validity of our findings. Finally, Section 9 concludes the paper and discusses implications for future research.
@@ 2. Related Work
In our literature review, we examine three key areas: 1) the user perspectives of Gen-AI apps, 2) user review analysis techniques, and 3) studies on prompt design and evaluation.
@@ 2.1. User Perspective of Gen-AI Apps
The user’s perspective on utilizing Gen-AI tools in education has been well studied (Golding et al., 2024; Kim et al., 2025b; Lee et al., 2024a). Shata and Hartley (Shata and Hartley, 2025) used surveys grounded in the Technology Acceptance Model (TAM) and Social Cognitive Theory (SCT) to examine faculty attitudes, identifying trust as the key driver of Gen-AI adoption. Kim et al. (Kim et al., 2025a) compared student and faculty perceptions on the use of Gen-AI in learning. Their study revealed that students and faculty used Gen-AI at similar rates, while students found Gen-AI tools easier to learn. Both faculty and students believed that Gen-AI will have more negative than positive effects on learning. Yuen and Schlote (Yuen and Schlote, 2024) surveyed adult learners to explore their views on the integration of AI in language learning apps. Their participants expressed a desire for in-app social features to assist in learning a language, and the authors suggested AI integration in the form of AI tutors that act as conversation partners.
@@
In the creative image domain, Tang et al. (Tang et al., 2024) surveyed users of Gen-AI image tools focusing on satisfaction, challenges, acceptance, perceptions, and applicability of Gen-AI image tools. They found that users most frequently rely on these tools during the early ideation stage rather than for producing final outputs. In support of that, Haase et al. (Haase et al., 2023) examined how Gen-AI images influence creative idea generation, and demonstrated that Gen-AI art effectively supports ideation. Oppenlaender et al. (Oppenlaender et al., 2023) focused more on the user’s understanding of technology, fears, and thoughts on the risks of Gen-AI image tools in their survey. Their participants voiced concerns over a societal risk, as people could use these tools to generate images that spread misinformation.
@@
In mental health, applications of AI include detection, diagnosis, treatment, and ongoing condition management (Sharma and Patel, 2024; Madhuri et al., 2025; S et al., 2025; Kheterpal and Gill, 2024). Our focus is on Gen-AI-driven treatment through conversational agents. Jin et al. (Jin et al., 2025) reviewed barriers to sustained engagement in mental health apps and suggested integrating AI in the form of AI agents that can detect emotional changes in users, build strong relationships with users, and promote timely intervention. Heinz et al. (Heinz et al., 2025) studied the efficacy of a mental health AI chatbot. They found that it significantly reduced users’ symptoms, maintained high engagement, and achieved user acceptance, with a reported therapeutic alliance comparable to that of human therapists.
Unlike prior work that relied on expert synthesis or self-reported perceptions, our study analyzes large-scale, real-world reviews from users actively engaging with Gen-AI apps. We offer a complementary perspective on actual usage and feedback. We also take a broad perspective by examining Gen-AI apps across all domains, without prioritizing any specific application area.

### v3

@@ Abstract.
The release of ChatGPT in 2022 triggered a rapid surge in generative artificial intelligence mobile apps (i.e., Gen-AI apps). Despite widespread adoption, little is known about how end users perceive and evaluate these Gen-AI functionalities in practice. In this work, we conduct a user-centered analysis of 676,066 automatically labeled reviews from 173 Gen-AI apps on the Google Play Store. We propose a structured four-phase framework, SARA (Selection, Acquisition, Refinement, and Analysis), which integrates and extends state-of-the-art techniques for large-scale review collection, filtering, and analysis using prompt-based LLMs. First, we empirically validate the reliability of LLMs for topic extraction and assignment, achieving 91% accuracy through five-shot prompting and LLM-based filtering of non-informative reviews. We then apply the framework to informative reviews to identify the ten most discussed topics (e.g., AI Performance, Content Quality, and Content Policy & Censorship) and analyze the key challenges and emerging opportunities. Finally, we examine how these topics evolve over time, offering insight into shifting user expectations and engagement patterns with Gen-AI apps. Based on our findings and observations, we present actionable implications for developers and researchers.
@@
$c$s: Software and its engineering Software maintenance tools
@@ 1. Introduction
Mobile applications (apps) are a crucial aspect of people’s lives, and their global data usage continues to rise each year (Rathod_Agal_2023). A major shift in the mobile app landscape occurred with the release of ChatGPT [^1] https://play.google.com/store/apps/details?id=com.openai.chatgpt&hl=en, which achieved 7.4 million downloads within its first ten days. This rapid adoption signaled strong consumer interest in generative artificial intelligence (Gen-AI) and sparked a surge of similar artificial intelligence (AI) powered apps across app marketplaces. In 2024, this trend contributed to a 112% year-over-year increase in downloads of AI chatbot apps (Ceci2025), such as Microsoft Copilot [^2] https://play.google.com/store/apps/details?id=com.microsoft.copilot and DeepSeek [^3] https://play.google.com/store/apps/details?id=com.deepseek.chat. Gen-AI refers to a class of AI that generates new and meaningful content, including text, images, audio and video (Feuerriegel_Hartmann_Janiesch_Zschech_2024; 10628027). Unlike traditional AI models, which rely on historical data to perform predictions and analytical tasks, Gen-AI models synthesize information from diverse sources to create original outputs (gozalobrizuela2023surveygenerativeaiapplications).
@@
Among the most prominent forms of Gen-AI are large language models (LLMs), i.e., deep learning models trained on massive text corpora to understand and generate human-like language (Hadi_Tashi_Qureshi_Shah_Muneer_Irfan_Zafar_Shaikh_Akhtar_Wu_et_al_2023). LLMs power a wide range of natural language processing (NLP) tasks, including text generation, summarization, translation, and dialogue. Developers are increasingly integrating Gen-AI models into mobile apps to enable novel capabilities, such as conversational interfaces, voice interaction, and content generation (Hau_Hassan_Zhou_2025). When embedded into apps, these models can deliver personalized and adaptive experiences by providing real-time, context-aware feedback that enhances user engagement and motivation (Yuen_Schlote_2024). A key enabler of this integration is the growing availability of Gen-AI-powered application programming interfaces (APIs). APIs serve as software intermediaries that enable mobile apps to access language understanding and generation functionalities without having to host or fine-tune models locally (Chen_Liu_Zhou_Zhao_Wang_Wang_Chen_Bissyandé_Klein_Li_2024).
@@
While the integration of Gen-AI into mobile apps offers clear benefits in functionality and user experience, it also introduces a range of technical and practical challenges. Implementing Gen-AI features requires consideration of API integration, prompt engineering, token limitation, inference cost, and the unpredictable model outputs (Chen_Gao_Chen_Zhang_Liu_2025). These issues can degrade the quality of Gen-AI-powered apps, leading to functional failures (e.g., crashes, incorrect behaviour, or suboptimal user interfaces), reduced efficiency (e.g., latency or excessive token consumption), and potential security vulnerabilities (Shao_Huang_Shen_Ma_Su_Wan_2024). Additionally, beyond implementation, quality assurance for Gen-AI-powered apps remains a concern. Unlike traditional software systems, Gen-AI components often lack formal specifications, making it challenging to evaluate correctness or performance in a deterministic way (Nahar_Kästner_Butler_Parnin_Zimmermann_Bird_2024). As a result, ensuring the reliability and robustness of Gen-AI-integrated mobile apps continues to be a significant challenge for developers.
@@
However, despite growing interest in integrating LLMs into mobile apps, little is known about the specific concerns and experiences of end-users. Existing research primarily focused on the challenges faced by developers. For example, Nahar et al. (Nahar_Kästner_Butler_Parnin_Zimmermann_Bird_2024) employed interviews and surveys to investigate how developers address Gen-AI integration issues, including latency, energy consumption, cost, and fairness. While informative, their work is limited to developer perspectives and does not consider how these technical challenges impact end-user experience. Chen et al. (Chen_Gao_Chen_Zhang_Liu_2025) analyzed posts from an OpenAI developer forum to uncover challenges in prompting, API use, and plugin development. This study offers insight into implementation bottlenecks but also neglects user-facing concerns. Similarly, Shao et al. (Shao_Huang_Shen_Ma_Su_Wan_2024) examined open-source mobile apps with Gen-AI integration and found that 77% experienced integration challenges, which impacted functionality, performance, or security issues. However, their analysis is limited to code-level integration problems and does not investigate whether users are aware of or affected by these issues.
@@
In contrast, our work examines the user-perceived impact of Gen-AI-enabled apps, using large-scale review data to understand how users perceive Gen-AI features in practice. Existing studies on user perspectives of Gen-AI apps tend to focus on specific domains, such as education (Golding_Lippert_Neuschatz_Salomon_Burke_2024; Kim_Yu_Detrick_Li_2025; Lee_Arnold_Srivastava_Plastow_Strelan_Ploeckl_Lekkas_Palmer_2024; Shata_Hartley_2025; Kim_Klopfer_Grohs_Eldardiry_Weichert_Cox_Pike_2025; Yuen_Schlote_2024), creative image generation (Tang_Zhang_Ciancia_Wang_2024; haase2023art; Oppenlaender_Silvennoinen_Paananen_Visuri_2023), or mental health (Jin_Kim_Han_2025; Heinz_Mackin_Trudeau_Bhattacharya_Wang_Banta_Jewett_Salzhauer_Griffin_Jacobson_2025). Our work takes a systematic view by analyzing user reviews across a wide range of Gen-AI tools and application areas. To our knowledge, this is the first study to systematically examine the full landscape of topics users discuss across diverse Gen-AI domains. We identify the technical pain points surfaced by users, in addition to emerging behavioural patterns, expectations, and sources of satisfaction or frustration. This user-centered perspective complements developer-focused studies, providing critical insights for improving the usability and effectiveness of Gen-AI mobile apps.
@@
To gain these insights, we analyze user reviews that offer valuable perspectives on the challenges, expectations, and satisfaction associated with Gen-AI apps. For example, Figure 1 illustrates representative reviews from three Gen-AI apps, highlighting both praise and frustration. Building on prior work in user review mining and LLM-based topic analysis (Maram_LLM_Cure; Chen_Lin_Hoi_Xiao_Zhang_2014), we propose a structured four-phase framework called Selection, Acquisition, Refinement, and Analysis (SARA). This framework enables the systematic collection, cleaning, and analysis of user reviews from Gen-AI apps, demonstrating the feasibility of large-scale, reproducible studies using LLMs.
First, we identify relevant Gen-AI apps (i.e., Selection) and collect user reviews from the Google Play Store (i.e., Acquisition). We then apply a data cleaning framework to enhance the quality of input data for LLM-based analysis (i.e., Refinement). Finally, we employ prompt-based LLM techniques to analyze the reviews (i.e., Analysis). Using this integrative framework, we automatically label and quantitatively analyze 676,066 user reviews from 173 Gen-AI apps using LLMs, complemented by qualitative analyses on statistically representative samples.
Before analyzing Gen-AI user reviews (RQ2–RQ3), we first evaluate the accuracy and reliability of our LLM-based framework (RQ1) to establish its validity for large-scale analysis. This ensures that the topic extraction and assignment process used in the subsequent analyses is accurate and trustworthy.
@@
In this RQ, we investigate how prompt configuration (e.g., using 0-shot, 3-shot, and 5-shot) and input quality (e.g., the proportion of informative reviews) affect the performance of LLMs in extracting topics from user reviews. Our results show that increasing the number of few-shot examples and filtering non-informative reviews significantly improve accuracy. The best-performing setup achieves an overall accuracy of 91%, confirming the reliability of our LLM framework for downstream analysis.
RQ2: What are the most prominent topics discussed in user reviews and in developer responses of Gen-AI apps?
To answer this RQ, we apply the LLM-based topic labeling framework from RQ1 to informative reviews across 173 Gen-AI apps. We analyze the top 10 topics in Gen-AI app reviews to understand user behaviour, key challenges and opportunities, and derive actionable insights for developers and researchers. While our primary focus is on user reviews, we also include a small supplementary examination of developer responses to better contextualize how developers engage with Gen-AI feedback. Our findings reveal that users tend to rate Gen-AI topics positively and expect high performance, often providing constructive feedback from a collaborative mindset. Notably, some users form emotional bonds with conversational agents and value personalized interactions. These insights have practical implications for developers aiming to improve the usability, engagement, and trustworthiness of Gen-AI features.
RQ3: What temporal trends emerge in user feedback on Gen-AI topics? In this RQ, we explore how user feedback on Gen-AI topics evolves as the technology matures. We cluster topics to identify evolutionary trends and support the analysis by qualitative review interpretation. Our analysis reveals that while the average rating for Gen-AI topics remains high, topic trends diverge: Content Quality ratings decline due to rising user expectations, AI Performance remains stable due to balancing effects, and Content Policy & Censorship ratings improve with better filtering mechanisms with time. These shifts reflect an increasingly critical and experienced user base that raises expectations for Gen-AI apps over time.
@@
  • We propose a structured four-phase framework, SARA, that builds upon established techniques in user review analysis and LLM-based topic modeling. Using SARA, we conduct a large-scale empirical study of 676,066 reviews from 173 Gen-AI apps on the Google Play Store to 1) identify the most frequently discussed Gen-AI-related topics, 2) uncover user expectations and pain points, and 3) analyze temporal trends to reveal how user perceptions evolve. While validated in the context of Gen-AI apps, SARA’s modular design makes it generalizable to analyzing user feedback in other mobile app domains and AI-integrated systems.
  • We evaluate the effectiveness of LLMs for topic extraction and assignment of user reviews, demonstrating that increasing the number of labeled examples and filtering out non-informative reviews significantly improves accuracy.
@@
  • We provide a replication package that includes the dataset of Gen-AI apps, three manually labeled samples for evaluating the filtering prompt, and the complete set of code we used to run all prompt-based experiments presented in this paper.
We organize the remainder of this paper as follows. Section 2 illustrates existing work on analyzing user reviews. Section 3 presents our methodology. Sections 46 describe the experiment design and the findings obtained for each RQ. Section 7 synthesizes our key findings and implications of our work. Section 8 discusses potential threats to the validity of our findings. Finally, Section 9 concludes the paper and discusses implications for future research.
@@ 2. Related Work
In our literature review, we examine two key areas: 1) the user perspectives of Gen-AI apps, and 2) user review analysis techniques.
@@ 2.1. User Perspective of Gen-AI Apps
The user’s perspective on utilizing Gen-AI tools in education has been well studied (Golding_Lippert_Neuschatz_Salomon_Burke_2024; Kim_Yu_Detrick_Li_2025; Lee_Arnold_Srivastava_Plastow_Strelan_Ploeckl_Lekkas_Palmer_2024). Shata and Hartley (Shata_Hartley_2025) used surveys grounded in the Technology Acceptance Model (TAM) and Social Cognitive Theory (SCT) to examine faculty attitudes, identifying trust as the key driver of Gen-AI adoption. Kim et al. (Kim_Klopfer_Grohs_Eldardiry_Weichert_Cox_Pike_2025) compared student and faculty perceptions on the use of Gen-AI in learning. Their study revealed that students and faculty used Gen-AI at similar rates, while students found Gen-AI tools easier to learn. Both faculty and students believed that Gen-AI will have more negative than positive effects on learning. Yuen and Schlote (Yuen_Schlote_2024) surveyed adult learners to explore their views on the integration of AI in language learning apps. Their participants expressed a desire for in-app social features to assist in learning a language, and the authors suggested AI integration in the form of AI tutors that act as conversation partners.
@@
In the creative image domain, Tang et al. (Tang_Zhang_Ciancia_Wang_2024) surveyed users of Gen-AI image tools focusing on satisfaction, challenges, acceptance, perceptions, and applicability of Gen-AI image tools. They found that users most frequently rely on these tools during the early ideation stage rather than for producing final outputs. In support of that, Haase et al. (haase2023art) examined how Gen-AI images influence creative idea generation, and demonstrated that Gen-AI art effectively supports ideation. Oppenlaender et al. (Oppenlaender_Silvennoinen_Paananen_Visuri_2023) focused more on the user’s understanding of technology, fears, and thoughts on the risks of Gen-AI image tools in their survey. Their participants voiced concerns over a societal risk, as people could use these tools to generate images that spread misinformation.
@@
In mental health, applications of AI include detection, diagnosis, treatment, and ongoing condition management (10842666; 10963316; 10985281; 10911418). Our focus is on Gen-AI-driven treatment through conversational agents. Jin et al. (Jin_Kim_Han_2025) reviewed barriers to sustained engagement in mental health apps and suggested integrating AI in the form of AI agents that can detect emotional changes in users, build strong relationships with users, and promote timely intervention. Heinz et al. (Heinz_Mackin_Trudeau_Bhattacharya_Wang_Banta_Jewett_Salzhauer_Griffin_Jacobson_2025) studied the efficacy of a mental health AI chatbot. They found that it significantly reduced users’ symptoms, maintained high engagement, and achieved user acceptance, with a reported therapeutic alliance comparable to that of human therapists.
Prior studies on user perceptions of Gen-AI apps have primarily relied on interviews and surveys to capture attitudes toward adoption, trust, and usefulness within specific domains such as education (Shata_Hartley_2025; Kim_Klopfer_Grohs_Eldardiry_Weichert_Cox_Pike_2025) and image generation (Tang_Zhang_Ciancia_Wang_2024; Oppenlaender_Silvennoinen_Paananen_Visuri_2023). These works offer valuable insight into how users conceptualize Gen-AI, but they often reflect processed opinions rather than immediate experiences. In contrast, our analysis of large-scale user reviews captures lived interactions, representing users’ spontaneous reactions written in context and often within moments of use. This perspective complements survey-based research by revealing practical concerns and responses as they naturally occur. For example, while prior studies in mental health (Jin_Kim_Han_2025; Heinz_Mackin_Trudeau_Bhattacharya_Wang_Banta_Jewett_Salzhauer_Griffin_Jacobson_2025) demonstrated that conversational AI can foster engagement and therapeutic rapport, our findings extend this understanding by showing that users explicitly value Gen-AI’s perceived neutrality, emotional support, and human-like realism. Furthermore, while earlier work examined single domains in isolation, our study provides a holistic view across multiple app categories, revealing shared topics and concerns that transcend context. By organizing insights by functional features rather than by domain, our findings highlight common usability, quality, and ethical challenges that characterize Gen-AI technologies more broadly.